S50, when a1 = 6 and d = 2/7
Answer by Cromlix(4381)

a(n) = a(1) +(n - 1)d
a(50) = 6 + (50 - 1)2/7
a(50) = 6 + 14
a(50) = 20
S(n) = [n(a(1)+ a(n)]/2
s(50) = [50(6 + 20)]/2
s(50) = 650
Hope this helps:-)